OpenStudy (anonymous):

Does anyone know how to do this problem? The height, s, in feet of a ball thrown straight up is given by the function s(t)=-16t^2+64t, where t is the time in seconds. Find the time it takes the ball to reach its maximum height. What is the maximum height?

OpenStudy (anonymous):

64 feet..There are two ways of doing this.. 1. Treat like a quadratic. Make -16t^2 + 64t + 0 = 0. The max will be when x=-b/2a. x=-64/(2*-32) = 2. When x = 2, y = 64ft, in this case x and y are t and s(t). The other way is with calculus. The max is reached when dy/dx (derivative) is 0. dy/dx of -16t^2 + 64t is -32t + 64. We set this equal to 0 and get t=2. Plug in and get 64. dy/dx is obtained by taking all terms and doing the following: take the exponent of the term, times the coefficient. That's your new coefficient. Then take the x term and make the exponent one less. Notice how the plane x becomes 1, because x is x to the 1st power. x to the zero power is 1.

OpenStudy (anonymous):

your vertex x=-b/2a x=t=-64/2(-16)=64/32=2sec max ht s=-16(2)^2+64(2)=64ft

OpenStudy (anonymous):

are you in calculus class tburn?

OpenStudy (anonymous):

No this is for intermediate algebra. I have a final tomorrow. This is one of the review problems. And I had no idea how to figure it out. Thank you so much for your help! You are awesom!

OpenStudy (anonymous):

ok god luck tburn for you exam